### v1

# Empathy Guidelines for Improving Practitioner Well-being & Software Engineering Practices

Abstract
Empathy is a powerful yet often overlooked element in software engineering (SE), supporting better teamwork, smoother communication, and effective decision-making. In our previous study, we identified a range of practitioner strategies for fostering empathy in SE contexts. Building on these insights, this paper introduces 17 actionable empathy guidelines designed to support practitioners, teams, and organisations. We also explore how these guidelines can be implemented in practice by examining real-world applications, challenges, and strategies to overcome them shared by software practitioners. To support adoption, we present a visual prioritisation framework that categorises the guidelines based on perceived importance, ease of implementation, and willingness to adopt. The findings offer practical and flexible suggestions for integrating empathy into everyday SE work, helping teams move from principles to sustainable action.
journal: IEEE Software
[…]
Human aspects play a significant role in software engineering (SE), shaping the interactions among software practitioners [1, 2]. One such human aspect is empathy, defined as “the ability to experience the affective and cognitive states of another person while maintaining a distinct sense of self” [3]. Empathy has been shown to support both the well-being of practitioners and the effectiveness of software development practices [4, 5, 6]. While empathy has been widely studied in fields such as medicine [7, 8], engineering [9, 10], and education [11, 12], it remains an underexplored topic in the context of SE [13].
Through interviews with 22 software practitioners [4], we identified both the benefits of empathy and the consequences of its absence. Empathy improved mental health, job satisfaction, collaboration, and technical outcomes such as code quality and project success. Its absence was linked to burnout, poor team cohesion, reduced productivity, higher turnover, declining code quality, and even the potential decline of the business. Practitioners also shared strategies for fostering empathy or addressing empathy-related challenges. Based on these insights, we developed 17 actionable guidelines to help individuals, teams, and organisations embed empathy in practice (Table 1). To develop these guidelines, we analysed the strategies from our earlier study [4] using the following criteria:
  • Participant support: We prioritised strategies that were frequently and consistently endorsed by participants, as these reflected strong practitioner backing.
  • Prevalence of causes of lack of empathy: We assessed how frequently each cause of a lack of empathy was reported by participants. Causes cited by a larger number of participants were considered more important, and strategies addressing these were given higher priority.
  • Real-world applicability: The strategies were analysed for their feasibility and practical implementation at different levels including practitioner, managerial, and organisational.
Table 1: Actionable empathy guidelines and their applicability at practitioner, managerial, and organisation levels
ID Guidelines Description Applicability of Guidelines
Practitioner Managerial Organisational
G1 Fostering strong relationships Strong relationships should be built among all stakeholders to $faCheckSquare \faCheckSquare$
among stakeholders promote empathy.
[…]
non-technical gap and non-technical stakeholders.
G3 Reducing friction among Interactions between different SE roles, such as developers and $faCheckSquare \faCheckSquare$
stakeholder groups testers, can lead to friction due to their contrasting responsibilities.
G4 Encouraging bi-directional Effective communication is essential for fostering empathy and $faCheckSquare \faCheckSquare$
communication building strong relationships.
[…]
business goals work with organisational goals.
[…]
to improve requirement stakeholders’ lack of empathy. Empathy in RE process helps
engineering developers better understand user needs, improving product quality.
[…]
solving technical issues, enhances solution quality by considering the
perspectives of all parties.
[…]
ceremonies helps address challenges & provide support during stand-ups, &
fosters a positive team culture in retrospectives and reviews.
[…]
G10 Creating a safe space Empathy plays a key role in fostering a judgment-free environment $faCheckSquare \faCheckSquare$
where practitioners can share concerns, enhancing well-being.
[…]
unexpected outcomes support & backup plans are vital to minimise project disruption.
G12 Flexibility in handling Empathy during personal crises fosters loyalty, while its absence $faCheckSquare \faCheckSquare$
human issues can lead to negative outcomes like resignations.
G13 Emphasising real-world Developers disconnected from end-users may overlook their work’s $faCheckSquare \faCheckSquare$
impact of developers’ work impact. Understanding this impact fosters empathy.
G14 Building an empathetic Company culture and leadership strongly influence practitioners’ $faCheckSquare \faCheckSquare$
team & company culture ability to express empathy, with leaders modelling empathy
encouraging others to do the same.
G15 Empathy education & training Integrating empathy into SE curricula & workplace training prepares $faCheckSquare \faCheckSquare$
professionals to apply empathy in daily practice.
[…]
(DEI) policies a safe, respectful, and inclusive workplace.
G17 Managing empathy fatigue Excessive empathy can cause fatigue, so setting boundaries and $faCheckSquare \faCheckSquare$
prioritising self-care is essential.
We conducted a large-scale survey [^1] Approved by Monash Human Research Ethics Committee. ERM Reference Number: 45708 with 103 software practitioners to explore the practical implementation of the empathy guidelines. Participants were recruited via Prolific [^2] https://www.prolific.com/, which provided access to a diverse, globally distributed, and pre-screened pool, avoiding the sampling bias of personal or professional networks. We applied targeted screening criteria including, participants were required to work in IT sector, have at least three years of experience, and regularly collaborate in teams. We included attention checks to ensure response quality and excluded responses that failed them. All open-ended questions were mandatory to encourage thoughtful feedback. Key demographic information are shown in Figure 1.
Practitioners were asked to evaluate the importance of each guideline in real-world industry settings, assess the practicality and ease of implementation of guidelines, and indicate their willingness to adopt these guidelines in everyday SE practices. In addition, we gathered qualitative insights on how practitioners would apply these guidelines in practice, the limitations they perceived, challenges they anticipated in adoption, and suggestions for improving the guidelines to enhance their applicability. Drawing on insights from this study, this article examines how the proposed empathy guidelines can be effectively implemented in practice.
Figure 1: Overview of Demographic Information of Large-Scale Survey Participants
Perceived Importance of Empathy Guidelines
Practitioners viewed empathy as fundamental to effective SE, enabling trust, psychological safety, and communication, particularly during agile practices and stakeholder engagement. It helped bridge gaps between technical and non-technical roles, aligning goals and improving outcomes, while its absence often led to misalignment and reduced quality. Empathy was also viewed as a motivator, connecting daily work to broader organisational goals, promoting user-centred thinking, and supporting well-being through empathetic leadership and flexible policies. Given that empathy may not come naturally to all, structured training was considered essential. While most supported the guidelines, some warned that over-empathising could lead to fatigue and distract from core responsibilities. Overall, embedding empathy into daily practice was seen to require strong leadership, supportive policies, and ongoing training.
Figure 2: A three-dimensional view of empathy guidelines’ readiness: importance (x-axis), ease of implementation (y-axis), and adoption willingness (bubble size) based on practitioner ratings.
Perceived Ease of Implementing Empathy Guidelines
Practitioners noted that many empathy guidelines were relatively easy to adopt, as they aligned well with existing SE practices. Guidelines such as fostering open communication, maintaining backup plans, and applying empathy during agile ceremonies were seen as natural extensions of standard workflows. These practices typically required little cross-team coordination, making them manageable within individual teams. In smaller or more cohesive teams, integrating empathy into agile routines or offering flexible work arrangements was considered especially straightforward. Overall, ease of implementation was closely tied to how well the guidelines fit within existing team processes and organisational culture. However, some guidelines were perceived as more difficult to apply due to structural, cultural, and resource-related challenges. These included limited managerial support for formal training, siloed communication in large organisations, and difficulty linking technical tasks to real-world impact. Tight schedules, cultural resistance, and a strong focus on deliverables over emotional engagement also hindered adoption. Sustained implementation was seen to require strong leadership and long-term commitment which were not always feasible amidst competing business demands.
Practitioner Willingness to Adopt Empathy Guidelines
Practitioners expressed strong interest in adopting the proposed guidelines to enhance empathy in their SE practice, citing a range of compelling reasons. Many noted that the guidelines support the creation of a safe, inclusive, and collaborative work environment, which they saw as fundamental to effective teamwork and cohesion. They emphasised that empathy enables developers to better understand and respond to the needs of users, stakeholders, and colleagues, resulting in more thoughtful and inclusive solutions. They noted that fostering open communication, trust, and strong team bonds not only enhances productivity but also contributes to a healthier and more resilient work culture. Several practitioners also highlighted empathy’s role in reducing stress, promoting well-being, and aligning team goals, which they viewed as important for long-term performance and satisfaction. The guidelines were also recognised as consistent with growing industry attention to psychological safety and emotional intelligence, with many participants reporting that they already applied similar principles in their work. Overall, the guidelines were considered practical and valuable for improving both project outcomes and the human experience of software development.
Figure 3: Practical approaches for implementing “Quick Wins” empathy guidelines (high importance, high ease) at the practitioner, managerial, and organisational levels.
Integrated Analysis of Importance, Ease, and Adoption Willingness
To synthesise these views of practitioners on the proposed empathy guidelines, we created a quadrant-based bubble chart that plots importance (x-axis) against ease of implementation (y-axis), with bubble size indicating practitioner willingness to adopt each guideline. Notably, some guidelines rated as less important or harder to implement still show large bubble sizes, indicating they may gain traction if supported under the right conditions. This visualisation enables categorisation of the guidelines into four strategic zones based on practical value and implementation feasibility, supporting prioritisation for real-world implementation.
Figure 4: Practical approaches for implementing “Nice to have” empathy guidelines (low importance, high ease) at the practitioner, managerial, and organisational levels.
Figure 5: Practical approaches for implementing “Strategic Investments” empathy guidelines (high importance, low ease) at the practitioner, managerial, and organisational levels.
  • Quick Wins: Top Right Quadrant (High Importance, High Ease)
    This quadrant includes the most promising guidelines, those that are both critical and easy to implement. Practitioners shared the practical ways of implementing our empathy guidelines in practice, as illustrated in Figure 3. Examples include G4 (Encouraging bi-directional communication), G5 (Ensuring transparency about business goals), G7 (Collaborative problem solving), and G10 (Creating a safe space). These guidelines not only received high importance ratings but were also perceived as relatively easy to adopt, with large bubble sizes reflecting strong practitioner willingness to implement them. In particular, G7 stands out with the highest adoption willingness in this group. These should be prioritised for immediate adoption in software teams.
  • Nice to have: Top Left Quadrant (Low Importance, High Ease)
    Guidelines in this area, such as G8 (Empathy during agile ceremonies), were considered relatively easy to implement but were rated as less critical. These can be seen as “low-hanging fruit”, simple enhancements that may still provide value when time and resources allow. These are low-risk, high-visibility actions that can still add value and support a more empathetic team culture. As illustrated in Figure 4, practitioners shared the practical ways of implementing these empathy guidelines in practice.
  • Strategic investments: Bottom Right Quadrant (High Importance, Low Ease)
    This quadrant includes guidelines perceived as very important but difficult to implement, such as G2 (Bridging the technical and non-technical gap) and G14 (Empathetic team and company culture). These are strategic investments, which require more effort, planning, or structural support but are essential for long-term cultural change. Practitioners may need organisational buy-in or policy-level support to effectively implement these recommendations. Practitioners shared the practical ways of implementing these empathy guidelines in practice, as illustrated in Figure 5.
    Figure 6: Practical approaches for implementing “Lowest priority” empathy guidelines (low importance, low ease) at the practitioner, managerial, and organisational levels.
  • Lowest priority: Bottom Left Quadrant (Low Importance, Low Ease)
    Guidelines here, such as G17 (Managing empathy fatigue) and G3 (Reducing friction), were rated both low in importance and difficult to implement. These are likely to be adopted more selectively or only in specific contexts where they align with team goals or organisational mandates. As illustrated in Figure 6, practitioners shared the practical ways of implementing these empathy guidelines in practice.
Limitations and Improvements
Based on practitioners’ insights, we identified key challenges in adopting the proposed empathy guidelines, along with practical strategies to overcome them, as shown in Figure 7.
Figure 7: Challenges in Applying Empathy Guidelines and Strategies to Overcome Them
Summary
Empathy plays a powerful role in making software teams more connected, effective, and user-focused. In this study, practitioners shared how empathy helps improve communication, build stronger relationships with stakeholders, and create healthier team environments. Many found the proposed guidelines easy to adopt, especially when they aligned with existing values or everyday practices. Others noted challenges such as time pressures, limited leadership support, and rigid organisational structures. Despite these challenges, they offered practical tips to embed empathy in daily work, through better communication, active listening, collaboration, and trust-building. To support real-world adoption, we also introduced an implementation prioritisation framework that helps teams decide which guidelines to focus on first, based on perceived importance, ease of implementation, and willingness to adopt. The key takeaway is that empathy is not merely a desirable addition, but a fundamental component of effective teamwork and high quality software development.
ACKNOWLEDGMENT
Gunatilake and Grundy are supported by ARC Laureate Fellowship FL190100035. We express our gratitude to all our participants for generously sharing their experiences.

### v2

@@
 Empathy is a powerful yet often overlooked element in software engineering (SE), supporting better teamwork, smoother communication, and effective decision-making.
-In our previous study, we identified a range of practitioner strategies for fostering empathy in SE contexts.
-Building on these insights, this paper introduces 17 actionable empathy guidelines designed to support practitioners, teams, and organisations.
+This paper introduces 17 actionable empathy guidelines designed to support practitioners, teams, and organisations.
 We also explore how these guidelines can be implemented in practice by examining real-world applications, challenges, and strategies to overcome them shared by software practitioners.
@@
 Empathy has been shown to support both the well-being of practitioners and the effectiveness of software development practices [9, 6, 3].
-While empathy has been widely studied in fields such as medicine [7, 8], engineering [9, 10], and education [11, 12], it remains an underexplored topic in the context of SE [13].
+While empathy has been widely studied in fields such as medicine [12, 16], engineering [14, 13], and education [1, 15], it remains an under-explored topic in the context of SE [10].
 Through interviews with 22 software practitioners [9], we identified both the benefits of empathy and the consequences of its absence.
-Empathy improved mental health, job satisfaction, collaboration, and technical outcomes such as code quality and project success.
-Its absence was linked to burnout, poor team cohesion, reduced productivity, higher turnover, declining code quality, and even the potential decline of the business.
-Practitioners also shared strategies for fostering empathy or addressing empathy-related challenges.
-Based on these insights, we developed 17 actionable guidelines to help individuals, teams, and organisations embed empathy in practice (Table 1).
-To develop these guidelines, we analysed the strategies from our earlier study [4] using the following criteria:
+Empathy was linked to improved mental health, job satisfaction, collaboration, and technical outcomes such as code quality and project success.
+In contrast, its absence was associated with burnout, poor team cohesion, reduced productivity, higher turnover, declining code quality, and even potential business decline.
+In this same study, practitioners also shared strategies for fostering empathy and addressing empathy-related challenges.
+Based on these insights, we developed 17 actionable guidelines to help individuals, teams, and organisations embed empathy in practice (Table 1) using the following criteria:
   • Participant support: We prioritised strategies that were frequently and consistently endorsed by participants, as these reflected strong practitioner backing.
@@
 Table 1: Actionable empathy guidelines and their applicability at practitioner, managerial, and organisation levels
-ID Guidelines Description Applicability of Guidelines
-Practitioner Managerial Organisational
-G1 Fostering strong relationships Strong relationships should be built among all stakeholders to $faCheckSquare \faCheckSquare$
-among stakeholders promote empathy.
+ID Guidelines Description Applicability in
+PL ML OL
 […]
-non-technical gap and non-technical stakeholders.
-G3 Reducing friction among Interactions between different SE roles, such as developers and $faCheckSquare \faCheckSquare$
-stakeholder groups testers, can lead to friction due to their contrasting responsibilities.
-G4 Encouraging bi-directional Effective communication is essential for fostering empathy and $faCheckSquare \faCheckSquare$
-communication building strong relationships.
+G2 Bridging technical (tech) and non-tech gap Empathy between tech and non-tech stakeholders enhances collaboration, and improves alignment between tech and business needs, leading to better SE outcomes.
+This guideline encourages assisting non-tech stakeholders better understand developers’ technical work, tailoring tech explanations to the audience’s understanding, and integrating cross-functional team members who understand both tech and non-tech aspects.
 […]
-business goals work with organisational goals.
+G4 Encouraging bi-directional communication Clear and effective two-way communication is essential for fostering empathy and building strong, collaborative relationships.
+Two-way communication establishes trust, minimises misunderstandings, and fosters mutual understanding, which are key elements in cultivating empathy.
 […]
-to improve requirement stakeholders’ lack of empathy. Empathy in RE process helps
-engineering developers better understand user needs, improving product quality.
+G6 Use an empathetic approach to improve RE Developers noted that unclear requirements often result from stakeholders’ lack of empathy.
+Empathy in RE process helps developers better understand user needs, improving product quality.
 […]
-solving technical issues, enhances solution quality by considering the
-perspectives of all parties.
+G8 Empathy during agile ceremonies Empathy during sprint planning supports realistic work allocation, helps address challenges & provide support during stand-ups, & fosters a positive team culture in retrospectives and reviews.
 […]
-ceremonies helps address challenges & provide support during stand-ups, &
-fosters a positive team culture in retrospectives and reviews.
+G10 Creating a safe space Fostering a safe environment where team members can share concerns without fear of judgement promotes well-being and positive team dynamics, enhancing mutual understanding and support.
+Empathy plays a key role in creating and maintaining this safe space.
 […]
-G10 Creating a safe space Empathy plays a key role in fostering a judgment-free environment $faCheckSquare \faCheckSquare$
-where practitioners can share concerns, enhancing well-being.
+G12 Flexibility in handling human issues Participants shared experiences where a lack of empathy during crisis situations led to negative outcomes, including resignations.
+In contrast, empathetic support allowed team members to handle emergencies and return with renewed loyalty.
+Demonstrating empathy and flexibility in these situations fosters team loyalty and contributes to project success.
 […]
-unexpected outcomes support & backup plans are vital to minimise project disruption.
-G12 Flexibility in handling Empathy during personal crises fosters loyalty, while its absence $faCheckSquare \faCheckSquare$
-human issues can lead to negative outcomes like resignations.
-G13 Emphasising real-world Developers disconnected from end-users may overlook their work’s $faCheckSquare \faCheckSquare$
-impact of developers’ work impact. Understanding this impact fosters empathy.
-G14 Building an empathetic Company culture and leadership strongly influence practitioners’ $faCheckSquare \faCheckSquare$
-team & company culture ability to express empathy, with leaders modelling empathy
-encouraging others to do the same.
-G15 Empathy education & training Integrating empathy into SE curricula & workplace training prepares $faCheckSquare \faCheckSquare$
-professionals to apply empathy in daily practice.
+G14 Building an empathetic team & company culture Company culture and leadership strongly influence practitioners’ ability to express empathy, with leaders modelling empathy encouraging others to do the same.
 […]
-(DEI) policies a safe, respectful, and inclusive workplace.
-G17 Managing empathy fatigue Excessive empathy can cause fatigue, so setting boundaries and $faCheckSquare \faCheckSquare$
-prioritising self-care is essential.
-We conducted a large-scale survey [^1] Approved by Monash Human Research Ethics Committee.
-ERM Reference Number: 45708 with 103 software practitioners to explore the practical implementation of the empathy guidelines.
-Participants were recruited via Prolific [^2] https://www.prolific.com/, which provided access to a diverse, globally distributed, and pre-screened pool, avoiding the sampling bias of personal or professional networks.
+G16 DEI policies Foster empathy through DEI policies that value diversity & promote a safe, respectful, and inclusive workplace.
+[…]
+PL: Practitioner Level, ML: Managerial Level, OL: Organisational Level, RE: Requirement Engineering, DEI: Diversity, Equity, and Inclusion
+In this paper, we introduce these guidelines for the first time and empirically evaluate them through a large-scale practitioner survey [8].
+We conducted a survey [^1] Approved by Monash Human Research Ethics Committee.
+ERM Reference Number: 45708 with software practitioners to explore the practical implementation of empathy guidelines.
+Participants were recruited via Prolific, [^2] https://www.prolific.com/ which provided access to a diverse, globally distributed, and pre-screened pool, avoiding the sampling bias of personal or professional networks.
 We applied targeted screening criteria including, participants were required to work in IT sector, have at least three years of experience, and regularly collaborate in teams.
@@
 All open-ended questions were mandatory to encourage thoughtful feedback.
-Key demographic information are shown in Figure 1.
+We collected a total of 125 survey responses.
+After removing responses that failed the attention-check criteria, 103 valid responses remained for analysis.
+Key demographic information is shown in Figure 1.
 Practitioners were asked to evaluate the importance of each guideline in real-world industry settings, assess the practicality and ease of implementation of guidelines, and indicate their willingness to adopt these guidelines in everyday SE practices.
@@
 Drawing on insights from this study, this article examines how the proposed empathy guidelines can be effectively implemented in practice.
+Recent work by Cerqueira et al.
+proposed a conceptual framework of empathy in SE based on a grey-literature analysis, offering a complementary perspective grounded in practitioners’ experiences [3].
+Their study identified workplace barriers such as toxic culture and an excessive technical focus, and integrated empathy practices with effects of empathy to present a holistic model.
+Our work differed in that we focused on empirically evaluating the practicality and adoption of developed empathy guidelines through a large-scale practitioner survey, and on identifying challenges to their implementation along with strategies to address these challenges.
+The qualitative data were analysed using thematic analysis [2].
+We conducted an inductive coding process to identify recurring themes related to how practitioners implemented the empathy guidelines in practice, the challenges they experienced, and the strategies they proposed to address these challenges.
 Figure 1: Overview of Demographic Information of Large-Scale Survey Participants
-Perceived Importance of Empathy Guidelines
-Practitioners viewed empathy as fundamental to effective SE, enabling trust, psychological safety, and communication, particularly during agile practices and stakeholder engagement.
+Perceived Importance
+Practitioners viewed empathy as fundamental to effective SE, enabling trust, psychological safety [4], and communication, particularly during agile practices and stakeholder engagement.
 It helped bridge gaps between technical and non-technical roles, aligning goals and improving outcomes, while its absence often led to misalignment and reduced quality.
@@
 Figure 2: A three-dimensional view of empathy guidelines’ readiness: importance (x-axis), ease of implementation (y-axis), and adoption willingness (bubble size) based on practitioner ratings.
-Perceived Ease of Implementing Empathy Guidelines
+Higher values on the x- and y-axes indicate higher perceived importance and greater ease of implementation, respectively.
+Larger bubbles indicate stronger willingness to adopt each guideline.
+This visualisation follows the scoring scheme used in the survey, where higher numerical values correspond to more favourable practitioner ratings, indicating greater ease rather than difficulty.
+All underlying numerical values used to construct the framework are available in our online Appendix(“Detailed Guidelines Analysis – Quant” sheet)[8].
+Perceived Ease of Implementation
 Practitioners noted that many empathy guidelines were relatively easy to adopt, as they aligned well with existing SE practices.
@@
 Sustained implementation was seen to require strong leadership and long-term commitment which were not always feasible amidst competing business demands.
-Practitioner Willingness to Adopt Empathy Guidelines
+Adoption Willingness
 Practitioners expressed strong interest in adopting the proposed guidelines to enhance empathy in their SE practice, citing a range of compelling reasons.
 Many noted that the guidelines support the creation of a safe, inclusive, and collaborative work environment, which they saw as fundamental to effective teamwork and cohesion.
-They emphasised that empathy enables developers to better understand and respond to the needs of users, stakeholders, and colleagues, resulting in more thoughtful and inclusive solutions.
-They noted that fostering open communication, trust, and strong team bonds not only enhances productivity but also contributes to a healthier and more resilient work culture.
+They emphasised that empathy enables developers to better understand and respond to others’ needs resulting in more inclusive solutions.
+They noted that fostering open communication, trust, and strong team bonds not only enhances productivity but also contributes to a healthier work culture.
 Several practitioners also highlighted empathy’s role in reducing stress, promoting well-being, and aligning team goals, which they viewed as important for long-term performance and satisfaction.
 The guidelines were also recognised as consistent with growing industry attention to psychological safety and emotional intelligence, with many participants reporting that they already applied similar principles in their work.
-Overall, the guidelines were considered practical and valuable for improving both project outcomes and the human experience of software development.
-Figure 3: Practical approaches for implementing “Quick Wins” empathy guidelines (high importance, high ease) at the practitioner, managerial, and organisational levels.
+Figure 3: Practical approaches for implementing “Quick Wins” empathy guidelines in green colour (high importance, high ease) and “Lowest priority” empathy guidelines in purple colour (low importance, low ease) at the practitioner, managerial, and organisational levels.
 Integrated Analysis of Importance, Ease, and Adoption Willingness
@@
 Practitioners shared the practical ways of implementing these empathy guidelines in practice, as illustrated in Figure 5.
-    Figure 6: Practical approaches for implementing “Lowest priority” empathy guidelines (low importance, low ease) at the practitioner, managerial, and organisational levels.
   • Lowest priority: Bottom Left Quadrant (Low Importance, Low Ease)
@@
 As illustrated in Figure 3, practitioners shared the practical ways of implementing these empathy guidelines in practice.
-Limitations and Improvements
+To support practical use, the prioritisation framework can guide teams in sequencing their adoption efforts.
+For example, a team experiencing communication breakdowns may begin with “quick wins” guidelines such as G4 (Encouraging bi-directional communication) or G7 (Collaborative problem solving), which offer immediate impact with minimal resource investment.
+Once these foundations are in place, the team may progress to “strategic investment” guidelines such as G2 (Bridging the technical and non-technical gap), which require broader organisational support.
+This phased approach enables teams to tailor adoption to their readiness and context while gradually building a more empathetic SE environment.
+Figure 6: Challenges in Applying Empathy Guidelines and Strategies to Overcome Them
+Overcoming Adoption Challenges
 Based on practitioners’ insights, we identified key challenges in adopting the proposed empathy guidelines, along with practical strategies to overcome them, as shown in Figure 6.
-Figure 7: Challenges in Applying Empathy Guidelines and Strategies to Overcome Them
+Practitioners highlighted several limitations that may hinder adoption, including time constraints, limited managerial support, and the cultural differences.
+In addition to these limitations, practitioners proposed a range of improvements to support successful adoption.
+These included integrating empathy-related practices into existing workflows, providing organisational backing through training and leadership endorsement, and promoting a culture that values open communication and psychological safety.
+Several participants emphasised that small, incremental changes such as regular check-ins, clearer expectations, and more intentional team interactions could make the guidelines easier to implement without imposing additional workload.
+Together, these challenges and strategies offer a clearer understanding of the conditions under which the guidelines are most feasible and the types of organisational and individual adjustments that could facilitate their uptake.
+Limitations
+Although Prolific offered global access, our sample still predominantly comprised participants from the Global North, which may have constrained the generalisability of our findings.
+In addition, cultural norms and organisational structures in different regions may influence how empathy guidelines are interpreted and applied.
+Future work should therefore recruit more participants from the Global South and examine these guidelines across varied cultural and organisational settings to strengthen representation and applicability.
 Summary
@@
 Many found the proposed guidelines easy to adopt, especially when they aligned with existing values or everyday practices.
-Others noted challenges such as time pressures, limited leadership support, and rigid organisational structures.
-Despite these challenges, they offered practical tips to embed empathy in daily work, through better communication, active listening, collaboration, and trust-building.
-To support real-world adoption, we also introduced an implementation prioritisation framework that helps teams decide which guidelines to focus on first, based on perceived importance, ease of implementation, and willingness to adopt.
+To support real-world adoption, we also introduced an implementation prioritisation framework that helps teams decide which guidelines to focus on first.
 The key takeaway is that empathy is not merely a desirable addition, but a fundamental component of effective teamwork and high quality software development.